OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is 15 and the last term is -4,199,040?

OpenStudy (anonymous):

so we know the general formula which is \[ar ^{n-1}\] where a is the first term and r is the multiplier. We are doing a geometric series so forget the n-1. Just go with n. \[ar ^{1} = 15\] We also know that the 8th term is \[ar ^{8} = -4, 199,040\] we can find r .\[\frac{ r ^{8} }{ r } = \frac{ -4,199,040 }{ 15 }\] solve for r and you get \[r ^{7} = -279,939\] take the 7th root and you get -6 so r = -6. Now to take the sum use the formula. \[Sn = a \frac{ [1-r ^{n}] }{[ 1-r]}\] So now find the sum by substitution.